Opened 10 years ago

Closed 9 years ago

Last modified 2 years ago

#267 closed flight processing (fixed)

GB08/10, flight day 268c, Denny Wood

Reported by: mark1 Owned by: knpa
Priority: alpha 4 high Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors: mark1

Description (last modified by emca)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB08_10-2009_268c_Denny_wood

Data arrived from ARSF via SATA disk L on the 7th October 2009

Scientific objective: see /users/rsg/arsf/arsf_data/2008/ARSF_Applications/GB_and_NL/GB08-10 Science Case.pdf

Priority: alpha-4-High

PI: Ross Hill

Sensors:

  • Eagle (14/12/2009)
  • Hawk (14/12/2009)
  • Leica LIDAR (26/04/2010)

Change History (26)

comment:1 Changed 10 years ago by mark1

Still awaiting flight logsheet.

DEM creation caused errors in Grass but still outputted a DEM. Check DEM is good when processing.

comment:2 Changed 10 years ago by mark1

Eagle and Hawk files in wrong project. Have moved the Eagle/Hawk files to boresight_2009-268a_little_riss.

Eagle/Hawk files from GB08_08-2009_268b_Wytham_woods have been moved into this project (during unpacking)

comment:3 Changed 10 years ago by mark1

Have renamed Eagle files 5-13 to have VNIR prefix

comment:4 Changed 10 years ago by mark1

  • Milestone set to 2009 Data processing completion

comment:5 Changed 10 years ago by mark1

logsheet arrived and unpacked

comment:6 Changed 10 years ago by mark1

have renamed hawk files from 268* to SWIR268*

comment:7 Changed 10 years ago by mggr

Re: #281 (Hawk frame rates are double what they should be), the following lines are affected and the .hdr file will need to be edited to halve the fps number:

uk/GB08_10-2009_268c_Denny_wood/hawk/SWIR268c-09-1.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_10-2009_268c_Denny_wood/hawk/SWIR268c-09-2.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_10-2009_268c_Denny_wood/hawk/SWIR268c-09-3.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_10-2009_268c_Denny_wood/hawk/SWIR268c-09-4.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_10-2009_268c_Denny_wood/hawk/SWIR268c-09-5.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_10-2009_268c_Denny_wood/hawk/SWIR268c-09-6.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_10-2009_268c_Denny_wood/hawk/SWIR268c-09-7.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_10-2009_268c_Denny_wood/hawk/SWIR268c-09-8.hdr: predicted length half recorded length, frame rate doubling problem?

comment:8 Changed 10 years ago by emca

  • Owner set to emca

comment:9 Changed 10 years ago by emca

Per flight line SCT values for Eagle and Hawk

Flight line Eagle Hawk
1 - 0.16
2 - 0.17
3 - 0.10
4 0.02 0.10
5 0.04 0.16
6 -0.01 0.18
7 0.04 0.10
8 0.07 0.11
9 0.04 -
10 0.11 -
11 0.01 -
12 0.05 -
13 0.06 -

comment:10 Changed 10 years ago by emca

Eagle lines 1, 2 and 3 are very short and are not included in the delivery (2 and 3 do not appear on the log sheet). Eagle line 7 has an incorrect number of lines in the header file and no dark frames. This line was processed using the dark frames from Eagle line 8.

Project is ready for delivery checking and delivery directory has been created at
~airborne/workspace/GB08_10-2009_268c_Denny_wood/delivery

comment:11 Changed 10 years ago by mark1

  • Other processors set to mark1

Starting delivery check

comment:12 Changed 10 years ago by mark1

Minor edits made to read me file, all data viewed and example commands tested (without dem). Data looks good.

Ready to deliver - data has not been rsynced back.

comment:13 Changed 10 years ago by emca

Eagle and Hawk data sent to PI 14/12/09

comment:14 Changed 9 years ago by emca

Starting LIDAR processing

comment:15 Changed 9 years ago by emca

LIDAR processing complete

Ready for delivery check at
~airborne/workspace/GB08_10-2009_268c_Denny_wood/delivery/20100422/GB08_10

comment:16 Changed 9 years ago by mark1

LIDAR delivery checked and all looks good.

Ready for delivery.

comment:17 Changed 9 years ago by emca

  • Component changed from Processing: general to Archiving
  • Description modified (diff)

LiDAR data dispatched to PI 26/04/2010

LiDAR and Hyperspectral rsync'd to
~arsf/arsf_data/2009/flight_data/uk/GB08_10-2009_268c_Denny_wood

Ready for archiving

comment:18 Changed 9 years ago by knpa

Starting to archive project.

comment:19 Changed 9 years ago by knpa

Missing applanix gps.

comment:20 Changed 9 years ago by knpa

Missing leica log files (no leica/logs dir).

comment:21 Changed 9 years ago by knpa

  • Owner changed from emca to knpa
  • Status changed from new to assigned

comment:22 Changed 9 years ago by knpa

No webcam data for this project.

comment:23 Changed 9 years ago by knpa

Finished preparing for archiving and have made available to NEODC for download.

comment:24 Changed 9 years ago by knpa

Backed up and archived by NEODC.

Have moved to archive disk.

comment:25 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleting workspace copy and closing ticket.

comment:26 Changed 2 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.